A surveyor wants to find the height of a building. At a point 111.0 from the base of the building he sights to the top of the building and finds the distance to be 191.0 ft. How high is the building (to the nearest tenth)?
Answer by Boreal(15235)

|
|
|==========
The hypotenuse is 191 and the adjacent side is 111; Pythagorean theorem.
111^2+x^2=191^2
36481=12321+x^2
24160=x^2
155.4 feet is the height.